Halwings grain company produces 1.5m dollars’ worth of grain with no inputs from other companies. General flour uses 1.5m dollars work of grain from Halwings to produce 3.2m dollars’ worth of flour. Finally hot and fresh baking buys 3.2 dollars’ worth of flour from General flour and uses it to produce 4m dollars’ worth of bread which it sells to its customers. Calculate the value of national output using the value added approach.



1
Expert's answer
2021-08-26T16:05:50-0400

Halwings value added = 1.5m – 0 = 1.5m

General flow value added = 3.2m – 1.5m = 1.7m

Hot and fresh value added = 4m – 3.2m = 0.8m

National value added (National Output) = 1.5m + 1.7m + 0.8m = 4m dollars